A copper penny dropped into a solution of nitric acid produces a mixture of nitrogen oxides. The following reaction describes the formation of NO, one of the products E cell for this reaction is 0.620 V 0.0350 M, Icu2 0.0400 M, and the partia at 208 K, when [H+] What is the value of E 0.100 M, [NO3 pressure of NO is 0.0018 atm? Number
Answer #1

Using Nernst equation to calculate Ecell at 208K

Ecell = E0cell - RT/nF * lnQ

as this reaction involves a 6-electron change (i.e. Cu ==> Cu2+ + 2e- is a 2-electron change, so 3 of those (3Cu ==> 3Cu2+) is a 6-electron change)

therefore, Ecell = 0.62- RT/nF * lnQ

= 0.62 - 0.059/6 log ((P NO)^2 [Cu2+]^3) / ([H+]^8 [NO3-]^2))

= 0.62 - 0.0098 log ((0.0018)^2 (0.04)^3) / (0.100)^8 (0.0350)^2)) = 0.62 - 0.0098 log 19.0 = 0.61 V